A horizontal force of 24 Newton acts on a body of mass 5 kg to move it's on a horizontal surface with acceleration of 3 m per second squared then the magnitude of the friction Force =
1
Expert's answer
2020-05-18T10:11:33-0400

given data:


"F_{horizontal} = 24 N"


"m = 5 Kg"


"a = 3 \frac {m} {sec^{2}}"


Net Horizontal Force


"F_{Net} = F_{horizontal} - F_{frictional}"


"F_{Net} = 24 - F_{frictional}"


Acceleration :


"a = \frac{F_{net}} {m}"


"a = \frac{24 - F_{frictional}} {m}"

"F_{frictional}=24- a\times m"


"F_{frictional}=24- a\times m"


"F_{frictional}=24- 3 \times 5"


"F_{frictional}= 9 N"